### v5

# Standalone 16-bit Neural Network Training: Missing Study for Hardware-Limited Deep Learning Practitioners

Abstract
With the increasing complexity of machine learning models, managing computational resources like memory and processing power has become a critical concern. Mixed precision techniques, which leverage different numerical precisions during model training and inference to optimize resource usage, have been widely adopted. However, access to hardware that supports lower precision formats (e.g., FP8 or FP4) remains limited, especially for practitioners with hardware constraints. For many with limited resources, the available options are restricted to using 32-bit, 16-bit, or a combination of the two. While it is commonly believed that 16-bit precision can achieve results comparable to full (32-bit) precision, this study is the first to systematically validate this assumption through both rigorous theoretical analysis and extensive empirical evaluation. Our theoretical formalization of floating-point errors and classification tolerance provides new insights into the conditions under which 16-bit precision can approximate 32-bit results. This study fills a critical gap, proving for the first time that standalone 16-bit precision neural networks match 32-bit and mixed-precision in accuracy while boosting computational speed. Given the widespread availability of 16-bit across GPUs, these findings are especially valuable for machine learning practitioners with limited hardware resources to make informed decisions.
Comprehensive data and detailed instructions for reproducing our results are publicly available at: https://github.com/YUNBLAK/standalone_16bits_nn.
Keywords: Deep Learning, Mixed Precision, Floating-Point, Hardware-Limited
1 Introduction
As machine learning models grow in complexity and size, the demand for computational resources such as memory and processing power has become a critical challenge. To address this, the mixed precision technique has been developed to balance resource consumption and computational efficiency [38, 53, 7, 55]. This approach strategically employs different numerical precision types during forward and backward propagation, weight storage, or inference. For instance, GPT-3 was trained with mixed precision of 32-bit and 16-bit formats [2]; Meta’s LLaMA 3 employs FP8 precision in both the inference and forward pass training phases [14, 44]. Additionally, much work involves quantizing pre-trained, full-precision models to lower precisions, such as int8 or even 1-bit, particularly for weight storage and inference [16, 56].
The trend toward mixed precision is closely linked to the development of lower-precision numerical formats. The ongoing ML arithmetic standard IEEE P3109, for example, has formalized seven 8-bit floating-point formats [1]; The Nvidia Blackwell chip is pioneering 4-bit precision; and research is even exploring the feasibility of ultra-low-bit neural networks for LLMs [37, 35]. Since these lower precision formats typically cannot function effectively on their own, mixed precision has become a must, now a dominant practice in machine learning.
Paradoxically, while mixed precision is designed to mitigate resource pressures in advanced computing, access to this technology often remains elusive for many machine learning practitioners facing hardware limitations. Table 1 shows some major GPUs in Nvidia and their hardware support for floating-point. Both FP32 and FP16 are supported across all GPUs listed. Support for bfloat16 begins with Nvidia’s Ampere architecture, whereas FP8 is supported starting from the Hopper architecture, and FP4 is available from the most recent Blackwell architecture [41, 42]. Note that, high-end, costly equipment such as Nvidia’s Hopper GPUs or BrainChip’s neuromorphic chips is not readily available to all.
One might argue that practitioners can have access to lower precision through software simulation and that some non-floating-point formats such as int8 or fixed-point do not require hardware acceleration [48, 9]. These options can be beneficial if used in weight storage or inference, but they are not practical due to the significant slowdowns in training time when relying on software simulation or without hardware acceleration. Therefore, for those practitioners constrained by hardware, the options for training models (to be differentiated from inference and storage) are typically limited to using 32-bit, 16-bit, or a combination of the two in mixed precision.
Among the three precisions available to hardware-constrained researchers, 16-bit or its mixed precision with 32-bit should be preferred theoretically, as they offer memory and performance advantages while being well-supported by modern hardware accelerators and frameworks. Yet, surprisingly, there’s not much information in current studies about the ability of the 16-bit formats to work by themselves. Despite the prevailing belief among machine learning practitioners regarding the benefits of 16-bit training, there has not been a definitive study specifically addressing its efficacy as a standalone solution. As is well known, there are two types of 16-bit precision: the IEEE binary16 and Google Brain’s bfloat16. Despite the growing interest in bfloat16, this work focuses on IEEE binary16 due to its wider accessibility, particularly for hardware-constrained practitioners. This decision is based on the fact that IEEE binary16 is more commonly available than bfloat16. Henceforth, we will refer to the 16-bit precision standardized by IEEE as 16-bit or FP16 interchangeably.
Table 1: Supported floating-point precisions across major NVIDIA GPU architectures. [41, 42]
Nvidia GPU Architectures Blackwell Hopper Ada Lovelace Ampere Turing Supported FP16, FP16, FP16, FP16, FP16, CUDA* FP32, FP64, FP32, FP64, FP32, FP64, FP32, FP64 FP32, FP64 Core precision bfloat16 bfloat16 bfloat16 Supported FP16, FP64, FP16, FP64, FP16, FP64, FP16, FP64, FP16 Tensor bfloat16, TF32, bfloat16, TF32 bfloat16, TF32 bfloat16, TF32 Core precision FP8, FP6, FP4 FP8 FP8
The missing study in standalone functionality of 16bit is likely due to two factors. Firstly, those not limited by hardware constraints are progressively shifting towards lower precision models such as 8-bit or 4-bit. Secondly, researchers experimenting with pure 16-bit neural networks observed numerical instability, which led to the proposal of mixed-precision techniques to address this issue [53]. So, 16-bit precision may have gained a reputation for being insufficient as a standalone solution. Indeed, direct attempts to train neural networks using 16-bit precision in frameworks such as TensorFlow or PyTorch typically fail. However, as will be discussed later, these issues are likely due to incorrect parameter settings. In particular, the default optimizer is Tensorflow and Pytorch is Adam, which updates weights based on a formula that relies a certain $\epsilon$ in the denominator (see Sect. 5.3 for details). In TensorFlow, the default value of $\epsilon$ is set at 1E-7, which shows to be too small for FP16 training. Meanwhile, PyTorch sets the default $\epsilon$ value at 1E-8, which effectively rounds to zero in FP16, as its smallest representable positive number is approximately 5.96E-8. Thus, we believe these two factors contribute to the situation that systematic research on standalone 16-bit training remains notably absent, a critical gap for hardware-constrained ML practitioners.
In this research, we aim to fill in this understudied area, standing from the perspective of resource-constrained ML practitioners to analyze standalone 16-bit training—both theoretically and practically. Here, we define "standalone" as employing 16-bit precision exclusively for training, weight storage, and inference, without the support of higher or lower precision formats in any phase. For benchmarking, we will compare standalone 16-bit training against 32-bit training and mixed precision methods. Our objective is to assess whether 16-bit training can independently serve as a feasible option. We believe that the results of this investigation could substantially improve the accessibility and utility of machine learning across a wider array of individuals and organizations.
Figure 1: Comparison of 16-bit floating-point neural networks with other precisions across different image classification architectures on CIFAR-10. Circle sizes indicate the number of parameters in each architecture. 16-bit, mixed precision, and 32-bit are represented by red, yellow, and blue colors, respectively. The x-axis denotes time, and the y-axis denotes accuracy.
Our work proceeds as follows. We start with a theoretical examination of the differences in accuracy between 16-bit and 32-bit. This theory includes formalizing concepts related to error tolerance and floating-point errors, proposing and proving a lemma that theoretically ensures that 16-bit can achieve comparable classification results to 32-bit under specific conditions. Preliminary observations support the hypothesis that 16-bit and 32-bit should perform similarly in handling classification problems. Then, through numerous comparisons against 32-bit and mixed-precision training, we highlight the relevance of standalone 16-bit training in image classification with CNN and ViT [13] architectures.
Fig. 1 illustrates some of our experimental results. All results were obtained using SGD optimizers rather than the default Adam optimizer, which likely contributed to the perception of 16-bit precision as insufficient for standalone use and, consequently, less studied (as mentioned above). To our knowledge, these results represent the first comprehensive evidence to suggest that standalone 16-bit floating-point neural networks can perform on par with their mixed-precision and 32-bit counterparts in classification. These findings should align with the expectations of many ML practitioners. However, it is essential to back such claims with scientific evidence to further the field. This work provides that missing piece of evidence through both theoretical and empirical perspectives.
In summary, this paper offers the following contributions:
  • Theoretical insights into the accuracy differences between 16-bit and 32-bit. Our theory provides a viable alternative to the perception noise in neural networks may act as regularization.
  • An extensive experimental study comparing 16-bit with 32-bit and mixed precision. We present a detailed empirical investigation to demonstrate the extent to which 16-bit floating-point can function as a standalone.
  • Identification of factors influencing 16-bit models. This paper further examines the factors that could influence the effective use of 16-bit models.
To facilitate reproducibility, we provide detailed descriptions of our experimental settings in the paper. We will release our code and evaluation results to the public upon publication.
2 Related Work
In the past decade, we have observed that scaling networks significantly improves their abilities to resolve complex tasks [31]. These increasingly large networks require extensive computations for their training and large-scale deployment. This trend raises several issues regarding accessibility, reproducibility, efficiency, environmental impact, and real-time execution. Given these critical challenges, numerous approaches have been proposed to address them and offer solutions to reduce energy and memory footprints while decreasing both running and training time. Quantization techniques reduce the precision of weights and activations to improve efficiency, as demonstrated by Jacob et al. [26], while methods like PACT [4] further optimize performance by learning clipping ranges. Pruning, as explored by Han et al. [20], removes redundant parameters to compress models, and Gale et al. [15] have shown its effectiveness in reducing resource usage without significant accuracy loss. Distillation, introduced by Hinton et al. [24], transfers knowledge from large models to smaller ones, and Touvron et al. [52] have adapted it for transformers. While these techniques are often viewed as primary solutions to these challenges, they are typically applied as post-processing steps before model deployment. Consequently, they do not reduce the training time and necessitate an additional step before deployment. In this work, we focus on the specific case of standalone IEEE 16-bit floating-point neural networks for both training and inference without any additional stages. Thus, this literature review will exclusively concentrate on works designed to accelerate both of these stages.
Low precision training pipeline. Training neural networks with low-bit precision while maintaining accuracy has been a critical focus in the literature. A key approach involves adapting the training pipeline to efficiently work with low-precision formats. For instance, early work [11] showed that using low-precision multiplications (e.g., 10-bit) can yield results comparable to 32-bit training, thanks to selective precision adjustments for different stages of computation. However, this approach can introduce challenges related to numerical stability and the risk of gradient underflow during training. Recent work such as LLM.int8() [12] introduces efficient 8-bit matrix multiplication techniques for transformers, significantly reducing computational cost while preserving accuracy. Nonetheless, this method requires careful handling of sensitive layers in the model, such as attention mechanisms, where quantization errors may arise.
At the core of low-precision approaches are two main types of numerical formats: floating-point and fixed-point. Floating-point numbers, commonly used in deep learning, can represent a wide range of values, which is important for complex tasks that require high precision. Fixed-point formats, on the other hand, use a fixed allocation of bits for integer and fractional parts, offering faster computation and lower memory usage, but at the cost of reduced precision [43]. This makes fixed-point formats more suitable for low-power, resource-constrained systems such as embedded devices and edge computing, although they may struggle with high-precision tasks that demand a wider dynamic range.
In the context of low-precision training, fixed-point strategies modify data representations while keeping the overall training pipeline unchanged. This simple and efficient approach is ideal for systems with limited computational resources. However, fixed-point formats often face challenges with range and precision, which can hinder their effectiveness in tasks that require finer numerical accuracy. Studies like [19], [3], and [36] have explored how fixed-point precision can lead to accuracy drops in neural networks and have proposed refinements to minimize these losses while maintaining efficiency. Nevertheless, these solutions may not always fully resolve the balance between efficiency and precision, especially in larger, more complex models where accuracy is more sensitive to the precision of weights and activations.
Moreover, integer quantization, a variant of fixed-point representation, has become an effective method for reducing computation costs. Several works, including [54], [10], and [26], have demonstrated the utility of integer quantization for neural network weights and activations. Despite its advantages in reducing computational overhead, integer quantization can lead to degradation in model accuracy, particularly in more complex models where maintaining sufficient precision is crucial. Recent advancements like SmoothQuant [56] and FlexPoint [33] dynamically adapt precision formats to further optimize memory and performance during training, but these approaches also require careful tuning to avoid accuracy loss. For example, studies such as Q8BERT [57] have applied 8-bit quantization to transformer models like BERT, demonstrating significant reductions in computational complexity. However, TernaryBERT [58], which applies 3-bit quantization to transformer models, reduces model size and energy consumption but presents challenges related to maintaining model accuracy, particularly in more detailed tasks, due to the lower bit-width.
Our research proposes a simple method that utilizes IEEE 16-bit floating-point operations for both training and inference of neural networks, without the need for complex integer quantization or dynamic application of various formats.
Mixed precision. The inception of a practical low-precision floating-point mechanism can be traced back to the work of Courbariaux et al. [8], who devised BinaryConnect, a method that reduces memory usage by binarizing weights during training while maintaining 32-bit precision for gradient updates, making it suitable for simpler models but potentially less effective for more complex ones. XNOR-Net [46] extended this by introducing binary operations into CNNs, further reducing computation but potentially sacrificing some accuracy in more detailed tasks, especially in vision tasks. Similarly, Jacob et al. [27] proposed an integer-only quantization method that applies quantization in the forward pass while retaining floating-point precision during backpropagation, balancing computational efficiency and model accuracy. However, the use of lower bit-widths may require careful calibration to avoid accuracy loss in more complex models.
Zhou et al. [59] introduced DoReFa-Net, which employs extremely low-bit representations (e.g., 2-bit activations and 4-bit weights) to minimize power consumption and memory usage, making it ideal for edge devices. However, the lower bit-widths require additional fine-tuning to maintain acceptable performance in complex tasks, and extreme quantization may result in significant accuracy loss if not calibrated correctly.
Micikevicius et al. [38] proposed a mixed precision training approach using 16-bit forward and backward propagations while maintaining 32-bit weights for updates. While this reduces memory usage, additional storage and conversion processes can sometimes offset the performance improvements. Wang et al. [53] proposed an 8-bit training method combined with 16-bit accumulations, significantly reducing computational load while preserving accuracy. Yet, as model complexity increases, ensuring sufficient precision can become challenging. Lastly, Koster et al. [29] revisited BF16, a 16-bit floating-point format with a broader dynamic range compared to FP16, making it more effective for deep learning tasks without significant accuracy loss. BF16 has gained widespread use in platforms like TPUs due to its ability to balance precision and performance in large-scale models.
Our approach differs by focusing solely on the use of IEEE 16-bit precision for both training and inference, without the need for mixed-precision adjustments or additional complexity, demonstrating that neural networks can achieve competitive accuracy using 16-bit precision alone.
Hardware support. Sharma et al. [50] introduced BitFusion, a dynamic hardware accelerator that adjusts bit-level operations at runtime to optimize resource use. This adaptive precision across neural network layers leads to significant energy savings and efficiency, making it ideal for low-precision tasks, though managing varying precision levels can introduce hardware design complexity. Similarly, fixed-point arithmetic, which reduces precision to cut computation and memory needs, is supported by hardware like FPGAs and ASICs [43]. In resource-constrained environments, such as embedded systems, fixed-point DSP processors are commonly used to accelerate low-precision operations [47].
Google’s Tensor Processing Unit (TPU) supports BF16, improving performance and memory efficiency in large-scale training [29]. Recently, NVIDIA’s Ampere and Hopper architectures expanded support for FP16, BF16, FP8, and FP4, further enhancing low-precision efficiency for both deep learning and large-scale models such as transformers [39]. While general-purpose GPUs support low-precision formats, specialized hardware such as BitFusion, TPUs, and fixed-point devices (e.g., FPGAs, DSPs) are essential for maximizing the benefits of low-precision computing, especially in large-scale training and edge applications. These technologies often require specific hardware environments optimized for energy efficiency and computational throughput.
Our work addresses this gap by rigorously investigating the efficacy of standalone IEEE 16-bit floating-point precision in various neural network architectures. Through both theoretical and experimental analysis, we provide evidence that standalone 16-bit precision can achieve comparable performance to 32-bit and mixed-precision approaches while offering significant improvements in computational efficiency. This study paves the way for more accessible, energy-efficient machine learning applications, especially for practitioners who may not have access to high-end hardware but still require efficient, low-precision solutions.
3 Theoretical Analysis
This section presents a theoretical comparison between standalone IEEE 16-bit and 32-bit floating-point deep neural networks. The objective is to study the reasons for any observed differences or similarities in their performance.
3.1 Background: Floating-Point Error
[…]
Rounding is necessary when representing real numbers that cannot be exactly represented. The rounding error of a real number $x$ at precision $p$ is defined as $|x-\circ_{p}(x)|$, where $\circ_{p}:\mathbb{R}\to\mathbf{F}_{p}$ is the rounding operation that maps $x$ to its nearest floating-point number in $\mathbf{F}_{p}$, namely $\circ_{p}(x)\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny def}}}}{{=}}}\mathrm{argmin}_{y\in\mathbf{F}_{p}}|x-y|$. (This definition of the rounding operation ignores the case with a tie for simplicity.)
Floating-point error includes rounding error and its propagation. For example, the floating-point code sin(0.1) goes through three approximations. First, 0.1 is rounded to the floating-point number $\circ_{p}(0.1)$ for some precision $p$. Then, the rounding error is propagated by the floating-point code sin proportional to its condition number [22]. Lastly, the output of the calculation is rounded again if an exact representation is not possible.
Floating-point error is usually measured in terms of absolute error or relative error. This paper uses the absolute error measurement, which quantifies the difference between two floating-point numbers $x$ and $y$ by a straightforward $|x-y|$.
3.2 Theory: Error vs. Tolerance
Notation.
A dataset $D\subseteq\mathcal{X}\times\{0,1,\ldots,n-1\}$ is a set of pairs of samples and labels. A classifier $M_{r}$ of precision $r\in\{16,32\}$, also known as classification model, is a function $\mathcal{X}\rightarrow([0,1]\cap\mathbf{F}_{r})^{n}$ corresponding to the 16-bit and 32-bit floating-point formats,
Definition 3.1.
Let $x\in\mathcal{X}$ be a sample input. Suppose $M_{16}$ and $M_{32}$ are 16-bit and 32-bit deep learning models trained by the same neural network architecture and hyper-parameters. We define the floating point error between the classifiers on $x$ as
$\delta(M_{32},M_{16},x)\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny def}}}}{{=}}}\max\{d|d\in|M_{32}(x)-M_{16}(x)|\},$ (1)
where $d\in|M_{32}(x)-M_{16}(x)|$ denote that $d$ is an element of the vector $|M_{32}(x)-M_{16}(x)|$.
Below, we investigate the degree to which the floating error affects the model accuracy. We first formalize the concept of classification tolerance and then prove a lemma, which provides a theoretical guarantee on the model accuracy of the 16-bit floating-point neural network with regard to its 32-bit counterpart.
Definition 3.2.
The classification tolerance of a classifier $M$ with respect to an sample $x\in\mathcal{X}$ is defined as the difference between the largest probability and the second-largest one:
$\Gamma(M,x)\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny def}}}}{{=}}}p_{0}-p_{1},$ (2)
where $p_{0}=\max\{p|p\in M(x)\}$, and $p_{1}=\max\{p|p\in M(x)\text{ and }p\neq p_{0}\}$. Here, $p\in M(x)$ denotes that $p$ is an element of the vector $M(x)$.
Lemma 3.3.
Let $\mathrm{class}(M_{r},x)$ denote the classification result of classifiers $M_{r}$ for $r\in\{16,32\}$ on a sample $x\in\mathcal{X}$. Namely, $\mathrm{class}(M_{r},x)\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny def}}}}{{=}}}\mathrm{argmax}_{i}\{p_{i}|p_{i}\in M_{r}(x)\}$. We have: If
$\Gamma(M_{32},x)\geq 2\delta(M_{32},M_{16},x),$ (3)
then $\mathrm{class}(M_{32},x)=\mathrm{class}(M_{16},x)$.
Proof.
Let $M_{32}(x)$ be $(p_{0},....p_{N-1})$. Without loss of generality we assume $p_{0}$ is the largest one in $\{p_{i}\}$ ($10\leq i\leq N-$ 1). We denote $\delta(M_{32},M_{16},x)$ by $\delta$. Following Eq. 3, we have
$\forall i\in\{1,...,N-1\},p_{0}-p_{i}\geq 2\delta.$ (4)
Let $M_{16}(x)$ be $(p^{\prime}_{0},....p^{\prime}_{N-1})$. Then for each $i\in\{1,...,N-1\}$, we have
[…]
$\geq p_{i}+\delta$ By Eq. 2 and Eq. 4
$\geq p_{i}^{\prime}$ By Eq. 1
Thus $p_{0}^{\prime}$ remains the largest in the elements of $M_{16}(x)$. ∎
Explanation of the lemma.
Suppose $M_{32}(x)=(p_{0},p_{1},\ldots,p_{n-1})$ for a classification problem with $n$ labels and let $p_{0}$ and $p_{1}$ be the largest and second largest probabilities in $\{p_{i}|i\in[1,n-1]\cup\mathbb{Z}\}$. Suppose $M_{16}(x)=(p_{0}^{\prime},p_{1}^{\prime},\ldots,p_{n-1}^{\prime})$. In the worst case (in terms of the difference between $M_{16}(x)$ and $M_{32}(x)$), $p_{0}^{\prime}$ can drop to $p_{0}-\delta(M_{32},M_{16},x)$, and $p_{1}^{\prime}$ can increase to $p_{1}+\delta(M_{32},M_{16},x)$. But as long as $p_{0}-\delta(M_{32},M_{16},x)>p_{1}+\delta(M_{32},M_{16},x)$, the two classifiers $M_{16}$ and $M_{32}$ must have the same classification result on $x$.
3.3 Observation: Standalone IEEE 16-bit Floating-Point DNN on MNIST
We validate Lemma 3.3 by comparing the performance of a standalone IEEE 16-bit floating-point deep neural network to a 32-bit one on MNIST. Our implementation is done using TensorFlow. The 16-bit implementation follows the same architecture as the 32-bit counterpart, except that all floating-point operations are performed using 16-bit precision.
Figure 2: Floating-point Errors vs. Classification Tolerance: Mean $\pm$ standard deviation
Figure 3: DNN Accuracies on MNIST Dataset: 32-bit vs. 16-bit floating-point
Fig. 3 presents the results of our experiments, showing the classification tolerance $\Gamma$ and floating-point error $\delta$. Our findings indicate that the mean floating-point error has a magnitude of approximately 1E-3 with a variance of up to 1E-4, while the classification tolerance is approximately 1E-1 with a variance of 1E-2. In other words, the average classification tolerance is about 100 times larger than the average floating-point error, suggesting that Eq. 3, namely $\Gamma>2\delta$, holds for most samples in the MNIST dataset. Additionally, we observed that in specific cases, floating-point errors can surpass the tolerance, reaching values as high as 9.95E-01 (presumably due to overflow/underflow). As a result, we anticipate that our 16-bit and 32-bit implementations will yield highly similar yet slightly different accuracy results.
Our experiments confirm this expectation. Fig. 3 shows the accuracy results on MNIST comparing the performance of the two implementations. We observe that the training accuracy of the 16-bit model is close to that of the 32-bit model, while the 16-bit model achieves slightly better accuracy in terms of testing accuracy.
This finding motivates us to investigate whether the similarity between the IEEE 16-bit and 32-bit models holds true for more complex neural networks. In theory, if a certain percentage, denoted by $a\%$, of samples in a dataset satisfy Eq. 3, Lemma 3.3 guarantees that the 32-bit and 16-bit models will have at least $a\%$ of their classification results in common. However, determining the exact $a\%$ poses a challenge. Nevertheless, we cautiously believe that for complex neural networks, a large majority of data satisfies Eq. 3. This belief stems from the fact that the loss function for classification problems, in the form of cross-entropy $-\Sigma\log(p_{i})$, guides the probabilities $p_{i}$ towards 1 during training, thereby resulting in a larger error tolerance $\Gamma$ compared to the relatively smaller $\delta$.
Remark.
Previous research has highlighted the ability of neural networks to tolerate noise and even act as a regularization method. However, our proposed analysis differs in two ways. Firstly, it is the first comparison of its kind that focuses on standalone IEEE 16-bit neural networks. Secondly, our analysis offers a unique quantification approach that distinguishes it from previous explanations related to regularization. Our approach examines floating-point errors and formalizes the concept of tolerance, allowing for its quantitative comparison with floating-point errors.
4 Experiments
We have conducted extensive experiments to address the question of whether IEEE 16-bit floating-point can function independently for image classification tasks. This section presents the results of our comparative study. We measured the wall-clock-time it took for the model to train for 100 epochs and test accuracy.
Result Summary
IEEE-16bit floating-point demonstrates standalone potential.
  • Ease of implementation: IEEE-16bit can be straightforwardly integrated, replacing 32-bit operations with 16-bit ones (Sect. 4.1). Generally, no fine-tuning is necessary (Sect. 5).
  • Accuracy: IEEE-16bit outperforms 32-bit in 3 out of the 12 examined architectures and exceeds Mixed precision in a 3 of them. On average, the accuracy of IEEE-16bit is slightly less than the accuracy of 32-bit and Mixed precision, 0.2% (Tab. 2). An everyday machine learning user may not notice distinction in accuracy (Fig. 4).
  • Runtime Efficiency: IEEE-16bit processes on average 1.6 times faster than 32bit and has a 1.3 times speed advantage over Mixed precision (Tab. 2).
Table 2: Performance metrics of standalone IEEE 16-bit, 32-bit, and mixed-precision methods [38] on CNN and Vision Transform architectures [13], using the CIFAR-10 dataset [30] without data augmentation and pre-trained models. All the experiments are repeated for 50 times, and we report the means and standard deviations.
Architectures Accuracy (%) Time (s) Accuracy Diff. Time Speedup (# par. million) FP16 FP32 MP FP16 FP32 MP FP16 - FP32 FP16 - MP FP32 / FP16 MP / FP16 AlexNet (2.09) [32] 76.0±$\pm{0.3}$ 75.8±$\pm{0.3}$ 75.9±$\pm{0.3}$ 96 174 150 0.2 0.1 1.8x 1.5x VGG-16 (33.76) [51] 83.7±$\pm{0.3}$ 83.9±$\pm{0.2}$ 83.8±$\pm{0.2}$ 377 857 455 -0.2 -0.1 2.2x 1.2x VGG-19 (38.36) [51] 83.8±$\pm{0.2}$ 83.9±$\pm{0.3}$ 83.9±$\pm{0.2}$ 416 937 492 -0.1 -0.1 2.2x 1.2x ResNet-32 (0.47) [21] 80.9±$\pm{0.3}$ 80.9±$\pm{0.4}$ 80.9±$\pm{0.3}$ 413 551 483 0.0 0.0 1.3x 1.2x ResNet-56 (0.86) [21] 81.6±$\pm{0.4}$ 81.4±$\pm{0.6}$ 81.5±$\pm{0.6}$ 677 905 795 0.2 0.1 1.3x 1.2x ResNet-110 (1.78) [21] 81.8±$\pm{0.4}$ 81.4±$\pm{0.5}$ 81.8±$\pm{0.5}$ 1256 1712 1486 0.4 0.0 1.3x 1.2x DenseNet121 (7.04) [25] 72.1±$\pm{0.3}$ 72.6±$\pm{0.4}$ 73.0±$\pm{0.3}$ 539 720 641 -0.5 -0.9 1.3x 1.2x DenseNet169 (12.65) [25] 71.7±$\pm{0.5}$ 72.3±$\pm{0.3}$ 72.1±$\pm{0.3}$ 724 966 812 -0.6 -0.4 1.3x 1.1x Xception (22.96) [5] 75.9±$\pm{0.4}$ 76.3±$\pm{0.3}$ 76.3±$\pm{0.4}$ 324 611 412 -0.4 -0.4 1.9x 1.3x MobileNetV2 (2.27) [49] 69.4±$\pm{1.1}$ 70.0±$\pm{1.2}$ 70.0±$\pm{1.1}$ 353 588 411 -0.6 -0.6 1.7x 1.2x VIT-8 (2.02) [13] 71.0±$\pm{0.3}$ 71.3±$\pm{0.3}$ 71.2±$\pm{0.3}$ 316 423 410 -0.3 -0.2 1.3x 1.2x VIT-12 (2.55) [13] 71.1±$\pm{0.3}$ 71.4±$\pm{0.3}$ 71.5±$\pm{0.3}$ 425 663 629 -0.3 -0.4 1.6x 1.5x Mean -0.2 -0.2 1.6x 1.3x
4.1 Baseline Precisions
We define three precision schemes denoted by FP16, FP32, and MP, respectively, which form the basis of our comparative study.
FP16:
This refers to the IEEE 16-bit floating-point precision. Our implementation is achieved from TensorFlow using tf.keras.backend.set_floatx(’float16’)[18]. Because operations involving both 16-bit and 32-bit can implicitly cast 16-bit values to 32-bit, we have cautiously examined the data types of the weights at various before, during, and post-training to ascertain they are indeed 16-bit and avoid unintentional mixture with 32-bit.
FP32:
This represents the conventional 32-bit floating point precision as defined by the IEEE standard. We utilize TensorFlow’s default implementation for this precision level.
MP:
The Mixed Precision scheme we refer to here is consistent with TensorFlow’s approach [18]. Our implementation is achieved from TensorFlow using tf.keras.mixed_precision.set_global_policy(’mixed_float16’). This methodology, initially introduced by Baidu and NVIDIA [38], is also supported in PyTorch, as highlighted in NVIDIA’s 2023 report [40].
4.2 Datasets and Architectures
For our image classification experiments, we employ the CIFAR-10 dataset [30], which consists of 60000 32x32 colour images in 10 classes. By default, this dataset is in 32-bit, but we convert it to IEEE 16-bit precision before training with FP16. Our experiments encompass 12 architectures: 10 established CNN architectures and the state-of-the-art Vision Transformer (ViT-8 and ViT-12) [13]. These CNN models include AlexNet [32], VGG16, VGG19 [51], ResNet-34, ResNet-56 [21], MobileNetV2 [49], DenseNet-121, DenseNet-169 [25], and Xception [5]. We don’t use pre-train weights or models for our experiments.
4.3 Measurement Metrics, Software and Hardware Settings
We carried out our experiments using 50 different random seeds, averaging both accuracy and training times to guarantee unbiased results. We carried out our experiments using 50 different random seeds, averaging both accuracy and training times(wall-clock-time) to guarantee unbiased results. Our tests spanned 100 epochs each, maintaining a consistent batch size of 256 and learning rate of 0.01 with the SGD optimizer [17] and momentum 0.9 [45]. We opted for the default settings, avoiding any parameter adjustments, and excluded data augmentation from our processes. All experiments have been conducted on a RTX4080 GPU and 40GB of RAM.
Figure 4: Comparative test accuracy over 100 epochs on CNNs and Vision Transformer (ViT) architectures utilizing IEEE 16-bit, 32-bit, and mixed precision.
4.4 Results
Figure 5: Boxplot of Test Accuracy: This figure illustrates the performance of CNN models and the Vision Transformer across three floating-point precisions: IEEE 16-bit, 32-bit, and mixed precision. Results from 50 random seeds are included to ensure unbiased representation. The white lines indicate the medians, while the white dots represent outliers.
Quantitative results are presented in Tab. 2, which outlines test accuracy and training duration with a side-by-side comparison of IEEE 16-bit Floating Point (FP16), 32-bit Floating Point (FP32), and mixed precision (MP). While it’s not obvious to single out an absolute winner, the 16-bit floating-point (FP16) clearly stands out in terms of computational speed.
One of the most striking observations is with the VGG-16 model, where FP16’s training time clocks in at a mere 377s, nearly half the time FP32 requires, 857s. This translates to a remarkable speedup of 2.2x. The efficiency of FP16 is further emphasized in Xception, with FP16 needing 324s compared to the substantial 611s that FP32 requires, again showcasing a 1.9x speedup.
When it comes to accuracy, FP16 performs competitively across most models. For example, in the AlexNet model, FP16 achieves an accuracy of 76.0%, slightly higher than FP32’s 75.8% and MP’s 75.9%. In VGG-16, FP16’s accuracy is 83.7%, which is slightly lower than FP32’s 83.9% and MP’s 83.8%. Similarly, FP16’s accuracy in ResNet-32 is 80.9%, matching FP32 and MP. In some models like DenseNet121, FP16’s accuracy is slightly lower than FP32 and MP by 0.5% and 0.9% respectively. On average, FP16’s accuracy is only 0.2% lower than FP32 and MP across the 12 examined architectures.
Despite the slight variations in accuracy, FP16 significantly improves computational training speed. On average, FP16 processes are 1.6 times faster than FP32 and 1.3 times faster than MP. This considerable increase in training speed makes FP16 an attractive option for practical applications where time efficiency is crucial.
Test accuracy comparison throughout the training process is depicted in Fig. 4. Neither FP16, FP32, nor mixed-precision emerges as a consistent frontrunner across all models. Fig. 5 offers a graphical summary of our outcomes across 50 random seeds showcased in a boxplot format. This encompasses metrics such as median, standard deviation, and outliers for the 12 architectures we evaluated. Based on median values, FP16 appears to slightly edge out the competition in VGG-16, ResNet-32, and ResNet-56. Concerning standard deviation, FP16 exhibits a smaller value across all architectures, with AlexNet being the sole exception. However, when assessing robustness, as indicated by outliers, FP16 demonstrates certain vulnerabilities, a shortcoming we recognize in Sect. 5.
In conclusion, our results indicate that utilizing IEEE 16-bit operations in CNN models and ViT significantly reduces run time without compromising accuracy, especially when compared to 32-bit and mixed precision.
5 Discussion
Although our results suggest that IEEE 16-bit floating-point can function effectively as a standalone, there are important caveats we would like to highlight.
Table 3: 16-bit (FP16) vs 32-bit (FP32) Precision Performance Comparison in Image Classification with Data Augmentation. MNIST [34] is used for DNN (Fully connected Network). CIFAR-10 Dataset [30] is used for ResNets [21] and VIT [13]. The results represent averages over experiments conducted with 4 random seeds and 300 epochs.
Model Test Accuracy Test Accuracy Acc Diff Speedup
FP16 FP32 (FP32 - FP16) (FP32 / FP16)
DNN 98.1±$\pm{0.2}$ 98.1±$\pm{0.3}$ 0.0 1.9x
ResNet-56 [21] 82.9±$\pm{0.3}$ 83.0±$\pm{0.3}$0.1 1.3x
ResNet-110 [21] 82.8±$\pm{0.3}$ 82.1±$\pm{0.4}$ 0.7 1.3x
VIT-8 [13] 74.2±$\pm{0.3}$ 74.5±$\pm{0.4}$0.3 1.4x
VIT-12 [13] 74.7±$\pm{0.3}$ 74.9±$\pm{0.3}$0.2 1.6x
5.1 Data augmentation
In our research, we initially did not include data augmentation to clearly demonstrate the effectiveness of 16-bit (FP16) precision models in straightforward image classification tasks on the MNIST dataset [34] and CIFAR-10 dataset [30]. Our main objective was to ascertain whether FP16 precision could be similar or exceed the performance of 32-bit (FP32) precision and mixed precision models. Following this, we introduced data augmentation to further test these models under more complex conditions. The experiment focuses on comparing the efficacy of 16-bit (FP16) and 32-bit (FP32) precision formats across DNN, two versions of the ResNet architecture — ResNet-56, ResNet-110 [21], and two ViT models [13].
In our experimental setup, we applied random left-right and up-down flips, adjusted the brightness by up to 10%, and rotated images within a range of -0.2 to 0.2 radians. The results, as shown in Table 3, compare FP16 and FP32 performances across different ResNet models, highlighting FP16’s consistent performance even with data augmentation. The data, averaged from experiments with four random seeds. This comprehensive approach, from non-augmented to augmented conditions, aimed to assess the viability of 16-bit precision in enhancing computational efficiency and accuracy across varied training environments.
5.2 Numerical Stability
Excluding outlier cases, the IEEE 16-bit performance is consistently on par with or better than both 32-bit and mixed precision in 4 of the 7 architectures we tested, as visualized in Figure 5. However, when considering these outliers, FP16 underperforms in terms of accuracy in 3 of the 7 architectures. This discrepancy is likely attributed to the known issues of overflow/underflow associated with the limited range of FP16. Delving into the reasons for this instability and potential countermeasures falls beyond the purview of this paper.
While such numerical instability might initially seem to deter the adoption of FP16 as a default choice, our results affirm that these anomalies are rare. Moreover, their effects can be counterbalanced by averaging results across multiple runs. This strategy can be particularly beneficial when working within strict memory constraints.
5.3 Hyperparameter Fine-tuning
As shown in Table 4, we conducted a performance analysis of the Adam optimizer [28] using float16 precision across different epsilon values. The results indicate that float16 does not perform optimally with the Adam optimizer when the epsilon parameter is set below $1\times 10^{-5}$. This suboptimal performance is particularly evident in the significant drop in test accuracy for epsilon values of $1\times 10^{-5}$ and lower.
Table 4: Performance analysis of the Adam optimizer was conducted using float16 precision format on CIFAR-10 [30]. It was observed that the float16 format did not exhibit optimal performance with the Adam optimizer [28] when the epsilon parameter was set below $1\times 10^{-5}$. Consequently, the Stochastic Gradient Descent (SGD) optimizer was employed as an alternative, which did not demonstrate any performance issues under these conditions.
FP16 ADAM
Epsilon 0.01 0.001 1e-4 1e-5 1e-6 1e-7 Time
ResNet-18 [21] 74.4 78.2 74.7 22.5 12.8 11.7 10.98s
ResNet-34 [21] 78.5 79.7 77.7 18.2 17.0 10.0 23.61s
To address this issue, we explored the use of the Stochastic Gradient Descent (SGD) optimizer, which did not exhibit the same performance problems under these conditions. Consequently, hyperparameter tuning for the 16-bit neural network was unnecessary when using the SGD optimizer.
$\text{RMSProp}:w_{t}\leftarrow w_{t-1}-\eta\frac{g_{t}}{\sqrt{v_{t}}+\epsilon}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (5)
$\text{ADAM}:w_{t}\leftarrow w_{t-1}-\eta\frac{\hat{m_{t}}}{\sqrt{\hat{v_{t}}}+\epsilon}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (6)
[…]
The analysis in Table 4 underscores the necessity of careful tuning of the epsilon parameter when using the Adam optimizer with float16 precision to avoid significant drops in performance.
5.4 Implementing 16-bit Batch Normalization
TensorFlow’s existing batch normalization layer doesn’t natively support 16-bit operations. Instead, it converts 16-bit inputs to 32-bit for calculations and then reverts back to 16-bit, introducing computational overhead due to the intermediate 32-bit processing. To meet our goal of evaluating standalone IEEE 16-bit neural networks, we implemented a dedicated 16-bit batch normalization layer. All our experimental results are based on this adapted layer, which fully supports 16-bit operations without requiring conversion to 32-bit.
6 Conclusion
In this study, we presented the first comprehensive theoretical and empirical analysis of standalone IEEE 16-bit floating-point neural networks. Despite the prevailing assumption that 16-bit floating-point neural networks can achieve results comparable to full (32-bit) precision, this study is the first to rigorously demonstrate this through both theoretical analysis and empirical evidence. This work fills a critical gap in the literature, providing rigorous proof that standalone IEEE 16-bit precision offers an effective and accessible option for reducing memory usage and improving computational speed, particularly for practitioners constrained by limited hardware resources.
The importance of this study is especially pronounced in resource-limited environments, where access to high-end hardware capable of supporting lower precision formats, such as FP8 or FP4, is often unavailable. Our findings show that standalone 16-bit precision can enable efficient machine learning training without the need for expensive or specialized hardware, making it a practical and valuable option for a wide range of users.
While our analysis highlights the benefits of 16-bit precision, we also identified certain limitations, particularly in relation to numerical stability in specific cases, which can affect performance. However, our theoretical formalization of floating-point errors and classification tolerance provides valuable insights into how 16-bit precision can approximate 32-bit results under the right conditions.
Overall, this work encourages broader adoption of standalone IEEE 16-bit floating-point neural networks, providing both theoretical and practical guidance. By addressing key challenges and demonstrating the feasibility of 16-bit precision in resource-limited settings, we believe this research opens up new possibilities for more accessible and efficient machine learning across diverse environments.
7 Statements and Declarations
The researcher claims no conflicts of interest. The datasets generated and/or analyzed during the current study, as well as the code for reproducing the results, are available on GitHub at https://github.com/YUNBLAK/standalone_16bits_nn.
Funding
This research was partially supported by the National Research Foundation of Korea (NRF) under Grant RS-2022-00165647. This work was also supported by the Innovative Human Resource Development for Local Intellectualization program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-RS-2023-00259678).